OpenStudy (anonymous):

the sum of a number and 1/8 of the number is 54. Find the number.

OpenStudy (aum):

Let the number be x x + x/8 = 54 Solve for x.

OpenStudy (anonymous):

wha t is the first step?

OpenStudy (aum):

Factor x out of the left hand side: x(1 + 1/8) = 54 Add 1 and 1/8. What is 1 + 1/8?

OpenStudy (anonymous):

would it be x(11/8) =54

OpenStudy (aum):

No. 1 + 1/8 = (8+1)/8 = 9/8 x * 9/8 = 54 Isolate x. multiply both sides by 8. divide both sides by 9.

OpenStudy (anonymous):

Im not understanding this yet. I am helping my brother with his math

OpenStudy (aum):

To add 1 and 1/8, find the least common denominator 1 can be written as 1/1 and to that add 1/8. The denominators are 1 and 8 and the least common denominator is 8. 1/1 + 1/8 = 8/8 + 1/8 = 9/8 9/8 * x = 54 Isolate x. Multiply both sides by 8: 9 * x = 54 * 8 Divide both sides by 9: x = 54 * 8 / 9 = ?